Sol Summary – April 10th

Crew 279 Sol Summary Report 10-04-2023
Sol: 8
Summary Title: Make hay while the sun shines
Author’s name: Antoine de Barquin, Commander & Agnès Dekeyser, Crew Executive Officer
Mission Status: Nominal
Sol Activity Summary:
Today was a hot day and we had it in our schedule. That’s why the EVA started an hour earlier than usual. The crew had to get up early to be ready to start at 8am. This allowed us not to be exposed to the sun when it was at its highest
Despite this precaution, taking radiation measurements for Thomas’s experiment around the base was no picnic. The hilly terrain behind the station and the already intense heat slowed down our progress. Finally the EVA was a success and we even had the time to take some drone pictures of the station.
In the meantime, the rest of the crew was preparing some good meals for the day. They tried to improve the bread recipe for the second time. Spoiler alert : it might have worked, they will discover the result at tonight’s dinner. They also tried a new recipe : cinnamon rolls! They were delicious but the dough was too hard. It seems that our martionauts have a new recipe to improve. Finally, they cooked some tuna pasta with tomato sauce with some vegetables that the whole crew ate with great pleasure.
The afternoon was filled with work such as taking pictures of the sun for some, turning some videos for others, or working on their own experiment for others. The time is going by extremely fast and our martionauts still have a lot to do but they all remain very motivated!
Look Ahead Plan:
Tomorrow, the crew will check the bacteria samples that were placed at the top of the North Ridge. They will also take this opportunity to perform some drone mapping as well as making some videos for their video report. The view that they will have of the MDRS will be perfect. The next EVA’s will be planned to continue Thomas and Augustin’s experiments meaning radioactivity measurements and drone mapping. In the MDRS, the crew will also continue their own experiment including data analysis and 3D printing.
Anomalies in work: nun-functional robotic observatory and the battery wires of Opportunity needs to be replaced
Weather: Sunny, high 24°C/75°F, low 2°C/36°F
Crew Physical Status: Nominal
EVA: See above
Reports to be filed: Journalist report, Greenhab report, Operations report, EVA report, EVA request
Support Requested: None

Operations Report – April 10th

Crew 279 Operations Report 10-04-2023

SOL: 8

Name of person filing report: Ioana Dimitrova

Non-nominal systems: robotic observatory, Opportunity
Notes on non-nominal systems:
Robotic observatory: awaiting for a solution from the manufacturer.
Opportunity Rover : see rover section

Spirit rover used: no
Hours: 215,9
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Opportunity rover used: no
Hours: 120,1
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Curiosity rover used: no
Hours: 228,5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
Perseverance rover used: no
Hours: 260,5
Beginning charge: 100%
Ending charge: 100%
Currently charging: yes
General notes on rovers: one of the battery wires on the driver side of Opportunity is corroded out and needs to be replaced.

Summary of Hab operations: Music influence on stress experiment was conducted. 3D print of Lunaria One, regolith collector. Radiation data inserted in 3D model.
WATER USE: 26,19
Water (static tank): 327,81
Static tank pipe heater (on or off): off
Static tank heater (On or off): off
Toilet tank emptied: no
Summary of internet: used from 14:00 to 21:00 for experiments and reports.

Summary of suits and radios: Radios charging nominal. Suits charging nominal.
Heater: Off
Supplemental light: Off
Harvest: 86g cherry tomatoes
Summary of ScienceDome operations: End of editing of mid-mission video.
Dual Split: Heat or AC, Off

Summary of RAM operations: none
Summary of any observatory issues: see above for robotic observatory. Solar observatory opened and used. See observatory report for more details.

Summary of health and safety issues: Slight fever and bloody noses. See HSO report.
Questions, concerns and requests to Mission Support:
We found a dead baby mouse behind the sink near the rear airlock…

EVA Report – April 10th

Crew 279 EVA Report 10-04-2023
EVA #9
Author: Thomas Stinglhamber, Crew Scientist
Purpose of EVA:

The EVA was the sequel to EVA #3 for Thomas’ experiment. He had to take radioactivity measurements near the base with his detector. He made punctual measurements of radioactivity (gamma emission) every 10 meters. The area was vast, so he aimed to do as many measurements as possible, targeting 100-150 measurements. He followed a snake-like route: North (250m)-West (10m)-South (250m)-West (10m)-North (250m), and so on. During EVA#3, he had taken measurements from the East of the Hab, and this time, he was covering the West part.
It had been difficult to make measurements with only one person for support during the last EVA, so he was helped by Antoine and Aglaé, who took notes and guided him in his tracing with the help of GPS.

Start time: 8:15 am
End time: 10:13 am
The day was likely to be very hot, so we began the EVA an hour earlier than usual to avoid exposure to the sun at its peak. The crew had to wake up early to prepare for an 8 am start. This was already the first difficulty. The second difficulty was the preparation of pancakes by Thomas and Antoine for the rest of the Crew… Fortunately the 2 difficulties were passed successfully.

Despite the precaution of getting up earlier to avoid the heat, taking radiation measurements for Thomas’s experiment around the base was difficult due to the hilly terrain and intense heat that began as early as 9:30 a.m. However, they persevered and managed to complete the EVA successfully. They made more than 140 measures which almost completed the planned map. In fact, they were so efficient that they even had time to capture drone footage of the station.

Destination: Mars Desert Research Station (MDRS)
Coordinates (use UTM WGS 84): 518200 E ,4251000 N
Participants: Thomas (EVA Leader, Scientist), Antoine (Commander), Aglaé (Journalist).
Road(s) and routes per MDRS Map: None
Mode of travel: By foot

GreenHab Report – April 10th

Crew 279 GreenHab Report 10-04-2023

GreenHab Officer: Augustin Tribolet

Environmental control: Open during day and heater during night

Average temperatures: 80.7 F

Hours of supplemental light: N/A

Daily water usage for crops: 16.9 gallons

Daily water usage for research and/or other purposes: N/A

Water in Blue Tank: 100.2 gallons

Time(s) of watering for crops: 06:44, 16:30, 20:45

Changes to crops: N/A

Narrative: The sun was intense today and it was very warm in the GreenHab. The door was left open all day. The plants will be watered again in the evening with 4.2 gallons (already accounted for in the water used today). The new planting of microgreens has started to germinate.

Harvest: N/A

Support/supplies needed: 86 g tomatoes

Astronomy Report – April 10th



Crew 279 Astronomer Report 10April2023
Dear Mr Detterline and MDRS support,

Today, I’ve also used the ZWO camera to capture the sun’s chromosphere and prominence. I think that I managed to capture more details in the sun chromosphere. However, there were no big prominences and I struggled a bit to blend the two exposures (chromosphere + prominence). Maybe I could use more time on this step.

Pictures settings :

Chromosphere : Gain=50, Exposure= 0.014, Gamma=1
Prominence : Gain= 210, Exposure=0.014, Gamma= 100
Flat : Gain=50, Exposure=0.014, Gamma=1

Concerning the green and blue frames of M51, tonight I will lunch again those to complete my observation.

I also attached a picture of my stargazing yesterday night.

Here is my report :

Astronomy Report

Name: le Bussy Gwenaël
Crew: 279
Date: 10/04/23

Robotic Telescope Requested (choose one) MDRS-14 (MLC-RCOS16) MDRS-WF

Objects to be Imaged this Evening: M51

Images submitted with this report: /

Problems Encountered: /


Solar Features Observed: Sun’s chromosphere and prominence

Images submitted with this report: Sun_23_04_10_chromosphere_and_prominence

Problems Encountered: blending exposure prominence+chromosphere
Best regards,

Journalist Report – April 10th

CREW 279 – Journalist Report – 10APR2023

Journalist Report : Aglaé Sacré, crew journalist
SOL 8: A sunny day

Today the day starts early. Considering the high temperatures, the EVA starts one hour in advance, at 8 am. For two hours, our three explorers continued to measure the radioactivity around the station for Thomas’ experiment.
The result of the addition of his data to Augustin’s mapping is impressive!
The other members of the team took the opportunity to sleep one hour more than usual!
During the night, we had a visit from little rodents.

As a little treat of the day, cinnamon rolls were made by our executive officer Agnès! A good way to rest with the heat today… and it’s not over yet.
This afternoon, we launched a 3D print for our GreenHab Officer experiment in collaboration with the company LunariaOne.

We also finalized our mid-mission video which is coming very soon!
Our astronomer could once again observe the sun which was very present today. And in the evening, he was able to observe the stars which were magnificent!

A great day again!

Copyright © The Mars Society. All rights reserved. | Main Site